### v1

# Part-Time Penalties and Heterogeneous Retirement Decisions

Part-Time Penalties and Heterogeneous Retirement Decisions
Kanta Ogawa [^1] A Ph.D. student at Graduate School of Economics, The University of Tokyo.
Contact: ogawa-kanta319@g.ecc.u-tokyo.ac.jp, kantaogawa319@gmail.com
I would like to express my sincere gratitude to Professor Kosuke Aoki at the University of Tokyo for his invaluable guidance as my master’s thesis advisor, and to Professor Sagiri Kitao at the National Graduate Institute for Policy Studies for her insightful advice. I also extend my thanks to the Recruit Works Institute, the Social Science Japan Data Archive, the Center for Social Research and Data Archives at the Institute of Social Science, The University of Tokyo, and the Panel Data Research Center at the Institute for Economic Studies, Keio University, for providing the data used in this research.
(March 23, 2025)
Older male workers exhibit diverse retirement behaviors across occupations and respond differently to policy changes, influenced significantly by the part-time penalty—wage reduction faced by part-time workers compared to their full-time counterparts. Many older individuals reduce their working hours, and in occupations with high part-time penalties, they tend to retire earlier, as observed in data from Japan and the United States. This study develops a general equilibrium model that incorporates occupational choices, endogenous labor supply, highlighting that the impact on the retirement decision is amplified by the presence of assets and pensions. I find that cutting employees’ pension benefits reduce aggregate labor supply in occupations with high part-time penalties, reducing overall welfare across the economy. In contrast, increasing income tax credits and exempting pension form income tax boost labor supply across all occupations and enhance welfare by raising disposable wages relative to the reservation wage. Reducing part-time penalties in high-penalty occupations also stimulate the labor supply in high-penalty occupations and improve long-term welfare.
Placeholder
Keywords: Retirement, Aging, Labor Supply, Occupational Choice
JEL Codes: E20, J26, J24
1 Introduction
Retirement timings vary significantly across occupations. While some professions tend to see workers retiring at sixty, others have mean retirement ages extending beyond seventy, as shown in Figure 1. Despite this striking heterogeneity, the factors driving such diverse retirement decisions remain unclear. Understanding these underlying factors is a requisite step to examine the impact of policy measures aimed at increasing the labor supply of older workers.
In aging economies, policymakers attempt to maintain the labor force size amidst demographic shifts. Policy reforms, such as raising the pension eligibility age and reducing pension benefits, have been introduced to encourage older workers to remain in the labor force. However, these reforms can yield strikingly heterogeneous effects across occupations. While some professions can respond positively by extending work participation, others may exhibit negligible or even negative responses, complicating the efficacy of such measures. Figure 2 illustrates the change in mean retirement age across occupations, comparing male cohorts born between 1918–1924 and those born between 1938–1944, using IPUMS-CPS data [Flood et al. [2024]]. The sample includes individuals who retired between the ages of 55 and 79, with data spanning from 1970 to 2024. While the earlier cohorts faced a normal retirement age of 65 under the pension system, the later cohorts experienced a gradual increase in the retirement age from 65 to 66 following the pension reforms enacted in the U.S. in 1983. Occupations are sorted by the magnitude of retirement age change, revealing substantial variation. While some occupations exhibit an increase in retirement age, a smaller subset of professions shows a decline. Although this trend may partially reflect broader time trends, it underscores the differential occupational responses to an extension of pension eligibility age.
This issue is particularly pressing as labor force is projected to decline further: working-age population in OECD countries is expected to shrink by 11% by 2062 compared to 2022 OECD [2023]. Given the prevalence of aging, which may amplify the variance in retirement decisions across occupations, it is plausible that policymakers should design retirement policies tailored to specific occupations.
This paper is the first to propose a framework to analyze how different occupations respond to policy changes, highlighting part-time penalties—wage reductions experienced by part-time workers compared to their full-time counterparts—as a key source of this heterogeneity. As illustrated in Figure 3, there is a negative correlation between part-time penalties and share of old workers: a proportion of male workers aged 60 and over among those aged 40 and over. The denominator includes both middle-aged and older workers to mitigate the influence of trends in occupational choice. Occupations with smaller part-time penalties tend to have higher rates of old workers as workers face less significant wage reductions when reducing working hours to spend more leisure time as they get older. In other words, in such occupations, workers are likely to retire earlier than in other occupations. This mechanism plays a crucial role in explaining the divergence in retirement timings across occupations.
As Blundell et al. [2016] provides a cursory overview of the general factors driving retirement discussed in the literature, much of the existing research has extensively examined retirement decisions in terms of health and social security systems. However, there is a notable oversight regarding the significant wage decreases faced by part-time older workers. While Rogerson and Wallenius [2013] argue that part-time penalties discourage older individuals from working part-time and often lead to permanent retirement without experiencing part-time roles, their analysis focuses on the general phenomenon. In contrast, my paper examines the varying degrees of part-time penalties across occupations and their role in shaping differences in retirement ages, incorporating features that enhance the understanding of this heterogeneity.
To begin with, this study accounts for the occupational heterogeneity of retirement decisions, building on the framework of Goldin [2014], which is developed to explain the narrowing gender wage gap. Following Jang and Yum [2022] and Erosa et al. [2022], which formalizes her concept within an equilibrium, the key differences between nonlinear and linear occupations are defined as part-time penalties, experience premiums, occupation-specific productivity, and age penalties. These characteristics characterizes the occupations in the model. While the classification primarily hinges on part-time penalties, the other three factors also play significant roles in explaining economic outcomes and worker behavior.
Furthermore, this paper first uncovers that part-time penalties play a more significant role than they may initially seem, as they interact with assets and pension benefits. The gist of the mechanism is as follows; as Goldin [2014] notes, occupations with high part-time penalties are typically high-skilled, offering greater compensation. Workers in these occupations tend to accumulate larger assets and expect more generous pension benefits, raising their reservation wage. Older individuals often experience increasing labor disutility due to declining health, the desire to spend more leisure time with their spouses, or the pursuit of hobbies, making them more inclined to reduce their working hours [^2] Another important consideration is highlighted by French and Jones [2012], which demonstrates that older individuals have higher labor elasticities compared to middle-aged workers.. In these circumstances, workers in high part-time penalty occupations face significant wage reductions [^3] Aaronson and French [2004] further demonstrates that transitioning to part-time jobs results in wage reductions for individuals in their early sixties. , making their potential earnings more likely to fall below their elevated reservation wage. Without switching to occupations with smaller part-time penalties, they are likely to exit the labor market permanently, as shown in Figure 10 in Appendix. Permanent exits are most frequent among those aged 55 to 79, followed by job switches within the same occupation category. Faced with significant part-time penalties444 Ameriks et al. [2020] examines a similar issue from a different angle, noting that the scarcity of jobs with flexible working conditions discourages older individuals from continuing to work., these workers are highly likely to choose permanent retirement. In contrast, workers in occupations with smaller part-time penalties typically continue working, as the wage reduction upon transitioning to part-time work is smaller.
Building on this concept, this research constructs a general equilibrium model of overlapping generations with endogenous labor supply, capturing both extensive and intensive margins with regard to labor, as well as occupational choices. Agents make decisions regarding consumption and savings, balancing the desire to leave a bequest or prepare for longevity while facing a survival shock each period. People unexpectedly become eligible to receive pension benefits at either age 60 or 65. This quantitative framework evaluates the impacts of policy reforms on different generations. In contrast to the literature555 While French [2005] and Fan et al. [2022] estimate life-cycle models to analyze retirement behavior, their approaches focus on a partial equilibrium. Similarly, Imrohoroğlu and Kitao [2012] demonstrates that social security reforms significantly affect the extensive and intensive margins of older individuals but do not incorporate occupational choices or part-time penalties. , my model highlights how changes in retirement behavior can significantly affect the welfare of other generations through shifts in labor supply, saving behavior, and prices, considering a general equilibrium effect.
To classify occupations and compute moments for quantitative analysis, this study utilizes the Japanese Panel Study of Employment Dynamics (JPSED). The JPSED covers more than 200 occupations and provides detailed personal information on each worker, including birth year, sex, education, work history, family status, and more. These rich variables enable precise regressions for classifying occupations. Additionally, the Japanese Household Panel Survey (JHPS/KHPS) supplements the analysis with asset data, which is not available in the JPSED. To ensure that similar retirement behaviors are observed in other countries, the IPUMS-CPS (Flood et al. [2024]), which provides data for the United States, is also employed.
Nonlinear and linear occupations are classified by regressing hourly wages on a quartic polynomial of working hours, controlling for factors such as age, birth year, family status, and others. The analysis reveals that nonlinear occupations tend to have a lower proportion of older workers, whereas linear occupations exhibit higher rates of older individuals.
The calibration analysis identifies part-time penalties as the primary source of nonlinearity in the model, followed by varying experience premiums across occupations, which reflect the increase in compensation from working additional periods. Moreover, counterfactual experiments are conducted to assess the impacts of policies aimed at increasing the labor supply of older individuals, and I find some interesting outcomes. The results indicate that eliminating the earnings test in pension rules666 Eliminating the earnings test has been shown to effectively increase the labor supply of older workers, particularly older males, in some countries (U.S.Blinder et al. [1980]; Friedberg [2000]; Song and Manchester [2007]; Haider and Loughran [2010]; Gelber et al. [2013], Canada: Baker and Benjamin [1999], U.K.: Disney and Smith [2002]). raises the intensive margin of older males by only $2.539\%$, while having minimal impact on welfare. This is consistent with empirical analysis in Japan[Shimizutani et al. [2008]], although research in other countries also demonstrate this policy change increases extensive margin of older males. Extending pension eligibility [^7] Numerous analyses also examine the extension of retirement ages across different countries and verify the resulting increase in the extensive margin of older workers (U.S.: Pingle [2006]; Mastrobuoni [2009], U.K.: Blundell and Emmerson [2003]; Cribb et al. [2013], Austria: Staubli and Zweimüller [2013]; Atalay and Barrett [2015], Switzerland: Hanel and Riphahn [2012]; Lalive and Staubli [2015]). and cutting pension benefits [^8] Several studies indicate that past pension reforms, including benefit reductions, have increased the labor force participation rate of older individuals (Anderson et al. [1999]; Gustman and Steinmeier [2009]; Blau and Goodstein [2010]; Brown [2013]). increases labor supply but slightly reduces output, leading to a welfare loss for all generations. When the pension eligibility age is extended, the capital supply decreases by $2.366\%$ in contrast to Imrohoroğlu and Kitao [2012] as workers adjust their retirement timing and experience flatter income profiles over time. In contrast, cutting pension benefits has varying effects across occupations: it reduces the labor supply in nonlinear occupations, as the working-age population becomes less motivated to increase working hours to boost future pension benefits. At the same time, it stimulates older workers in nonlinear occupations to remain in the workforce, highlighting the heterogeneous occupational responses to such policy changes. In this case, the former effect outweighs the latter.
I propose several unconventional policies to increase the labor supply of older workers while positively impacting overall welfare. For instance, reducing the part-time penalties in nonlinear occupations increases labor supply in the nonlinear occupations by $2.247\%$ and enhances long-term welfare in terms of consumption equivalence. Policies such as raising tax credits and exempting pensions from income tax are also effective in boosting labor supply across both types of occupations, increasing output, and improving welfare.
Section 2 discusses empirical facts, and section 3 elaborates on the model. Section 4 presents the calibration results. Section 5 details the counterfactual experiments, and section 6 concludes this paper.
Figure 1: Mean Retirement Timings of Each Occupation: Males, 2015-2019
Figure 2: Change in Mean Retirement Age by Occupation: Males Born 1918–1924 vs. 1938–1944 in US
Figure 3: Part-time Penalty and Share of Old Workers: Males, 2015-2019
88footnotetext: The part-time penalty is rescaled such that the average hourly wage is standardized to one.
”Share of Old Workers” represents the proportion of male workers aged 60 and older among those aged 40 and older in each occupation.
2 Empirical Analysis
2.1 Data
The Japanese Panel Study of Employment Dynamics (JPSED), collected by the Recruit Works Institute and released by the University of Tokyo, provides data for classifying occupations and calibrating the model. Covering over 100 individual attributes from 2015 to 2022, it includes approximately 50,000 annual observations. For classification, all valid observations are used with controls for time and sex. For calibration, the analysis focuses on males aged 25–104 from 2015 to 2019, avoiding the COVID-19 pandemic’s heterogeneous impact on retirement decisions.
Household asset data is supplemented with the Japanese Household Panel Survey (JHPS/KHPS), with KHPS starting in 2004 and JHPS in 2009. Asset moments are calculated using data from 2012 to 2019, including financial and housing assets, while pre-2015 data increases observations for older individuals. KHPS and JHPS provide approximately 3,000 and 2,500 annual observations, respectively, ensuring alignment with model calibration.
For robustness checks, IPUMS-CPS data from 2009 to 2024 is used to examine whether a similar pattern is observed in the U.S. The dataset, which includes approximately 130,000 to 220,000 individuals, provides detailed information on personal attributes and work-related characteristics. Occupations are classified based on this data, allowing for a comprehensive analysis of labor market trends and retirement behavior.
2.2 Occupational Classification Strategy
I classify occupations using regression analysis, addressing limitations in methods used by Erosa et al. [2022] and Jang and Yum [2022], which categorize occupations based on male working hours. This approach is unsuitable for Japan, where most males aged 25–59 work full-time, leading to unintended results. For example, truck drivers and barbers, with longer working hours, often exhibit linear wage-hour relationships, while researchers and IT engineers, with shorter hours, show nonlinear patterns. To resolve this, I adopt a regression-based method, controlling for factors like age, sex, and family status, using data from both sexes aged 25–79 to capture more part-time workers.
I conduct the classification in the following procedure:
  1. 1.
    For each of the more than 200 occupations, I regress hourly wages on a quartic polynomial of weekly working hours, controlling for age (as a polynomial), sex, education, marital status, child status, residential area, and time-fixed effects. Occupations with fewer than 200 observations are excluded.
  2. 2.
    I calculate the residualized hourly wage difference between individuals working 10 hours per week and those working 40 hours per week. I define this wage difference standarized by the average hourly wage of males aged 25-79 as a part-time penalty.
  3. 3.
    I classify the top 50 % of occupations with the largest part-time penalties as nonlinear, and the remainder as linear [^9] Nearly all occupations classified as nonlinear exhibit positive part-time penalties, whereas linear occupations typically have negligible or negative penalties..
For example, IT engineers, researchers, pharmaceutical sales representatives, and banking sales representatives are classified as nonlinear occupations. In contrast, construction workers, cooks, and character and CG designers are classified as linear occupations.
Figure 4 illustrates the change in hourly wage, normalized to the hourly wage at 10 hours per week, between nonlinear and linear occupations. The figure shows that hourly wages increase more rapidly with working hours rise in nonlinear occupations, which aligns with the original concept of these occupational categories. Unlike Goldin [2014], who assumes no wage changes in linear occupations, I allow for minor wage increases.
The data supports the hypothesis that workers in nonlinear occupations face high part-time penalties, leading to earlier retirement compared to those in linear occupations. As shown in Figure 5, the share of nonlinear occupations among working males steadily declines after age 60, while the share of linear occupations rises. Table 1 shows that non-working rates increase sharply from 4.74% (ages 25–59) to 77.10% (ages 70–79). Before age 60, most workers in both occupation types work full-time. After age 60, the share of workers decreases more sharply in nonlinear occupations, while linear occupations see a smaller decline as workers continue with reduced hours.
Figure 4: Hourly Wage Change over Working Hours in Japan(nonlinear vs. linear), 2015-2019
Figure 5: Change of Conditional Occupation Share over Age in Japan (nonlinear vs. linear), 2015-2019
[…]
Age
25-59 60-69 70-79
Occupation
Annual Working Hours(h) Nonlinear Linear All Nonlinear Linear All Nonlinear Linear All
00 4.743 30.56 77.10
$(0,1000)$ 0.2373 0.5963 0.8336 1.499 3.278 4.777 1.411 3.632 5.043
$[1000,1500)$ 0.5223 1.325 1.847 2.662 5.871 8.533 1.439 3.699 5.138
$[1500,2000)$ 5.194 4.753 9.767 7.268 8.567 15.84 1.653 3.198 4.851
$[2000,2500)$ 28.61 25.58 54.19 13.42 18.82 32.24 2.071 4.140 6.211
$2500\leq$ 12.71 15.74 28.45 2.505 5.549 8.054 0.4214 1.236 1.657
$0<$ 47.27 47.99 95.26 27.35 42.09 69.44 6.574 15.91 22.48
99footnotetext: The table provides an unconditional proportion of workers in each category. For example, between the ages of 25 and 59, 4.743% of individuals do not work, while 95.26% are employed. 47.27% of agents work in nonlinear occupations, and 0.2373% of individuals work less than 1,000 hours per year in nonlinear occupations .
3 Model
In this section, I present the details of the model. I integrate the occupational choice model developed by Jang and Yum [2022] into the retirement decision framework to explain the heterogeneity in the proportion of older workers among overall employees across different occupations.
3.1 Demographics
A continuum of males is born each year at age 25 (t=1) and lives until age 104 (t=80). A fraction of these individuals begins receiving pension benefits at age 60 (t=36), while the remainder start at age 65 (t=41). Pension eligibility is determined at age 60 as a random shock, remaining unknown until that point. All agents retire from the labor force by age 80 (t=56). Each agent faces a survival shock in every period. Upon death, their bequests are evenly distributed among the remaining survivors.
3.2 Preferences
Each agent has preferences over consumption, labor supply, and assets which are denoted by $c$, $h$., and $a$, respectively. They decide whether to work and, if working, select occupations. Utility is derived from consumption, while disutility arises from labor supply, which consists of two components: a fixed cost of working, $\xi$, and labor disutility, $\Phi$, which starts to increase at $R+1$. All agents have this utility function:
$u(c,h;\phi,t,a_{0})=\frac{c^{1-\sigma}}{1-\sigma}-\Phi(t)\frac{h^{1+\frac{1}{\gamma}}}{1+\frac{1}{\gamma}}-\xi\mathbbm{1}\{h>0\}$
where the coefficient of labor disutility is given by $\Phi_{i}(t)=\phi_{i}+\kappa(t-R)\mathbbm{1}\{t>R\}$ [^10] I set R=35, which denotes 59 years old in real terms, which means that the labor disutility starts to increase when the agent turns 60 years old., and the fixed cost of working is expressed by $\xi$.
Agents also have a bequest motive, and all bequests are equally distributed among surviving agents. This model focuses on accidental bequests [^11] I refer to De Nardi [2004] to formulate this bequest motive. and excludes the inheritance of earnings ability and inter-vivo transfers. The utility derived from leaving a bequest is modeled as:
$\mu(a^{\prime})=\mu_{1}\bigg{(}1+\frac{b(a^{\prime})}{\mu_{2}}\bigg{)}^{1-\sigma}$
where $b(a^{\prime})$ denotes the after-tax bequest. $\mu_{1}$ represents the agent’s concern for leaving bequests, and $\mu_{2}$ indicates the extent to which bequests are considered luxury goods.
3.3 Pension
Pension benefits are composed of two terms: the national pension, $\underline{b}$, which is distributed equally to all agents, and employees’ pension insurance, which is based on the agent’s past earnings. The mean of the agent’s past labor earnings is updated each period using the following equation,subject to an upper bound on labor earnings, $\bar{e}$, when calculating pension benefits. Until age 70, $e$ is updated according to the rules of employees’ pension insurance.
$ssb(e)$ $=\underline{b}+\rho e$
$e$ $=\frac{e_{-1}\times(t-1)+\min\{\lambda,\bar{e}\}}{t}$
, where $\lambda$ denotes today’s post-tax labor income.
3.4 Efficiency Labor
Each agent is compensated by firms based on their efficiency labor, which depends on several factors. A working agent provides an effective labor supply to the representative firm, and wages are paid per unit of effective labor. If an agent of age $t$ and experience $x$ works in occupation $j$ for hours $h$ per week, their income is given by:
$w_{j}\underbrace{m_{j}(\eta_{1})g_{j}(h)z_{j}(x,j_{-1},t)f_{j}(t)}_{\text{Effective Labor Supply}}$
First, the worker selects an occupation based on the occupation-specific productivity, $m(\eta_{1})$ [^12] NL and L refer to nonlinear and linear occupations, respectively.. A worker draws $\eta_{1}$, an idiosyncratic value, at birth, which remains constant throughout their lifetime. If a worker is well-suited for occupation $j$, they typically remain in the same job until retirement. The occupation-specific productivity is defined as:
$m(\eta_{j})=\begin{cases}e^{\eta_{1}}&(j=\text{NL})\\ 1&(j=\text{L})\end{cases}$
Second, the part-time penalty is governed by the function $g_{j}(h)$, which depends on labor supply. As an agent works more hours, their productivity increases, with the degree of this increase varying across occupations. In general, productivity rises more significantly in nonlinear occupations compared to linear ones. The following functional form is assumed when calibrating the parameters:
$g_{j}(h)=h^{1+\theta_{j}}$
where $\theta_{j}>-1$.
Moreover, productivity also depends partly on experience within the occupation. A worker accumulates one unit of experience for each period of work [^13] The process of accumulating experience does not require full-time work, as only 8.9% of males between 25 and 59 work less than 35 hours per week in the data. This full-time work constraint would be necessary if the focus were on analyzing the gender wage gap..
$z_{j}(x,j_{-1},t)=\begin{cases}1+\underbrace{\Omega_{j}\min\{x,\bar{x}_{j}\}}_{\begin{subarray}{c}\text{Experience}\\ \text{Premium}\end{subarray}}&(j=j_{-1})\\ 1&(j\neq j_{-1})\end{cases}$
Experience is updated according to the rule: $x=x_{-1}+1$ if $j=j_{-1}$ and $x=0$ if $j\neq j_{-1}$, where $x_{-1}$ represents prior experience. Each unit of an experience adds $\Omega_{j}$ units of an experience premium until reaching the upper bound, $\bar{x}_{j}$ [^14] From the data, I set $x_{NL}=35$ and $x_{L}=29$, which, respectively, denote 59 and 53 years old., which varies across occupations. As long as the worker remains in the same occupation, they continue to accumulate experience. However, if the worker switches to a different occupation, their experience resets, starting from $x=0$.
[…]
$f_{j}(t)=\exp\Big({-\pi_{j}(\min\{t,\bar{R}_{j}\}-R)\mathcal{I}_{t>R}}\Big{missing})$
3.5 Household Problem
Using these features, I construct a household problem that accounts for both the extensive and intensive margins of labor supply, allowing workers to choose their occupations. Productivity increases with age until experience reaches $\bar{x}_{j}$ after which it begins to decline at age $R+1$. Pension eligibility starts unexpectedly for some agents at age 60, while others begin receiving benefits at age 65, with amount based on past earnings. All agents retire by age 80 and live up to 104, facing survival shocks.
First, I describe the problem faced by agents between the ages of 25-79. Each period, agents decide on their consumption, next-period assets, labor supply, and, if working, their occupations in the current period. In this problem, a worker decides whether to work in period $t$, given the state $(a,x_{-1},j_{-1},\phi,\eta_{1},t,e_{-1},p)$. Here, $a$ denotes the current assets, and $x_{-1}$ represents the years of experience in the current occupation. The variable $j_{-1}$ determines the experience premium in combination with $x_{-1}$, because if a worker switches occupations, their experience is reset to zero, and they must start from scratch. The variable $\phi$ represents an idiosyncratic coefficient for labor disutility function, $\Phi$, and $\eta_{1}$ is a parameter in the nonlinear occupation-specific productivity, which determines the worker’s suitability for each occupation. The variable $t$ represents the worker’s age, and $e_{-1}$ is the mean of the worker’s past earnings, which determines the amount of pension benefits. The variable $p$ represents pension eligibility: if $p=1$, the agent is eligible for pension benefits. All younger agents are ineligible, meaning $p=0$. A fraction of the population starts receiving a pension at age sixty, while others become eligible at age sixty-five. Agents do not know their exact pension eligibility age until they turn sixty and begin receiving pension benefits if eligible.
The post-income tax function, $\mathcal{Y}(\cdot)$, takes three inputs: financial before-tax income, labor income, and pension benefits. I replicate Japan’s 2019 tax system, as there were no significant tax reforms during the period used for calibration.
People make these decisions simultaneously every period, solving the following maximization problem:
$V^{Y}(a,x_{-1},j_{-1},\phi,\eta_{1},t,e_{-1},p)\\ =max\Big{\{}N(a,\phi,\eta_{1},t,e_{-1},p),W(a,x_{-1},j_{-1},\phi,\eta_{1},t,e_{-1},p)\Big{\}}$
Here, $N(\cdot)$ and $W(\cdot)$ correspond to the value functions of not working and the value of working, respectively. The decision to work or not is represented by $n\in\{NW,W\}$, where NW indicates not working and W indicates working.
Next, if working, an agent selects an occupation. $J_{j}$ is the value of working in occupation $j$, where $j=NL$ and $j=L$ represent nonlinear and linear occupations, respectively. For convenience, I also denote $j=NW$ to represent a non-worker.
$W(a,x_{-1},j_{-1},\phi,\eta_{1},t,e_{-1},p)=max\Big{\{}J_{NL}(a,x_{-1},j_{-1},\phi,\eta_{1},t,e_{-1},p),J_{L}(a,x_{-1},j_{-1},\phi,\eta_{1},t,e_{-1},p)\Big{\}}$
The value function of occupation $j$ is clearly defined by:
$J_{j}(a,x_{-1},j_{-1},\phi,\eta_{1},t,e_{-1},p)$
$=max_{\begin{subarray}{c}c,a^{\prime}\geq 0,h\in[0,1]\end{subarray}}\bigg{\{}u(c,h)+\Big{(}1-S(t)\Big{)}\mu(a^{\prime})+\beta S(t)\mathbb{E}\Big{[}V^{Y}(a^{\prime},x,j,\phi,\eta_{1},t+1,e,p^{\prime})\bigg{]}\Big{\}}$
subject to:
$c+a^{\prime}=a+Tr+B+\mathcal{Y}\Big{(}ra,w_{j}m_{j}(\eta_{1})g_{j}(h)z_{j}(x,j_{-1},t)f_{j}(t),\mathcal{I}_{p=1}ssb(e)\Big{)}$
, where $\beta$ denotes a discounted factor.
In the budget constraint, $Tr$ and $B$ denote the public lump-sum transfer and bequest from the deceased, respectively. They survive to the next period with a probability $S(t)$ and die with a probability $1-S(t)$, leaving a bequest. The agent earns income from assets, labor, and, if eligible, a pension. They allocate disposable income—after paying social security contributions and taxes on labor and financial income—towards consumption and asset accumulation.
Alternatively, if not working, the agent faces the value of not working:
$N(a,\phi,\eta_{1},t,e_{-1},p)=max_{c,a\geq 0}\bigg{\{}u(c,0)+\Big{(}1-S(t)\Big{)}\mu(a^{\prime})+\beta S(t)\mathbb{E}\Big{[}V^{Y}(a^{\prime},0,0,\phi,\eta_{1},t+1,e,p^{\prime})\Big{]}\bigg{\}}$
subject to:
$c+a^{\prime}=a+Tr+B(t)+\mathcal{Y}\Big{(}ra,0,\mathcal{I}_{p=1}ssb(e)\Big{)}$
Lastly, after age 80, agents no longer work and rely solely on interest from assets and pension benefits, which continue to depend on their past earnings.
$V^{O}(a,t,e_{-1},p^{\prime})=max_{c,a\geq 0}\bigg{\{}u(c,0)+\Big{(}1-S(t)\Big{)}\mu(a^{\prime})+\beta S(t)\mathbb{E}\Big{[}V_{O}(a^{\prime},t+1,e,p^{\prime})\Big{]}\bigg{\}}$
subject to:
$c+a^{\prime}=a+Tr+\mathcal{Y}\Big{(}ra,0,\mathcal{I}_{p=1}ssb(e)\Big{)}$
3.6 Representative Firm
I elaborate on the settings of the production sector. The representative firm demands capital, nonlinear labor, and linear labor.
Nonlinear labor and linear labor for occupations are denoted by $L_{1}$ and $L_{2}$, respectively, while K represents the firm’s capital demand.
$max_{L_{1},L_{2},K}Y-w_{1}L_{1}-w_{2}L_{2}-(r+\delta)K$
,where
$Y$ $=$ $AK^{\alpha}(L)^{1-\alpha}$
$L$ $=$ $\Big{[}\nu L_{1}^{\frac{\psi-1}{\psi}}+(1-\nu)L_{2}^{\frac{\psi-1}{\psi}}\Big{]}^{\frac{\psi}{\psi-1}}$
The first-order conditions are as following:
$r$ $=\alpha\bigg{(}\frac{L}{K}\bigg{)}^{1-\alpha}-\delta$
$w_{1}$ $=(1-\alpha)\bigg{(}\frac{K}{L}\bigg{)}^{\alpha}\nu\bigg{(}\frac{L}{L_{1}}\bigg{)}^{\frac{1}{\psi}}$
$w_{2}$ $=(1-\alpha)\bigg{(}\frac{K}{L}\bigg{)}^{\alpha}(1-\nu)\bigg{(}\frac{L}{L_{2}}\bigg{)}^{\frac{1}{\psi}}$
3.7 Stationary Competitive Equilibrium
I assume that the initial asset $a_{0}$ is zero, and I define individual states, $s_{Y}$ and $s_{O}$, as $s_{Y}\equiv(a,x_{-1},j_{-1},\phi,\eta_{1},t,e_{-1},p)\in\textbf{S}^{Y}$ and $s_{O}\equiv(a,t,e_{-1},p)\in\textbf{S}^{O}$, respectively, where the state spaces, $\textbf{S}^{Y}$ and $\textbf{S}^{O}$, are $\textbf{S}^{Y}\equiv\textbf{A}\times\textbf{X}\times\textbf{J}\times\Phi\times\textbf{$\eta$}\times\textbf{T}\times\textbf{E}\times\textbf{P}$ and $\textbf{S}^{O}\equiv\textbf{A}\times\textbf{T}\times\textbf{E}\times\textbf{P}$.
[…]
  1. 1.
    Given ($r,w_{1},w_{2}$), policy functions $\{c_{Y}(s_{Y}),a_{Y}^{\prime}(s_{Y}),h_{Y}(s_{Y}))\}_{s_{Y}\in\textbf{S}^{Y}}$, $\{c_{O}(s_{O}),a_{O}^{\prime}(s_{O})\}_{s_{O}\in\textbf{S}^{O}}$ and $\{n(s_{Y}),j(s_{Y})\}_{s_{Y}\in\textbf{S}^{Y}}$ and value functions $\{V^{Y}(s_{Y}),N(s_{Y}),W(s_{Y})\}_{s_{Y}\in\textbf{S}^{Y}}$ and $\{V^{O}(s_{O})\}_{s_{O}\in\textbf{S}^{O}}$ solve the household problem.
    The working decision rules are determined by
    Extensive $\text{Extensive Margin}:n(s_{Y})$ $=$ $argmax\{N(s_{Y}),W(s_{Y})\}$
    Occupational $\text{Occupational Choice}:j(s_{Y})$ $=$ $argmax\{J_{NL}(s_{Y}),J_{L}(s_{Y})\}$
    where $n(s_{Y})\in\{NW,W\}$ and $j(s_{Y})\in\{NL,L\}$.
  2. 2.
    Given ($r,w_{1},w_{2}$), $K^{D}$, $L_{1}^{D}$ and $L_{2}^{D}$ solve the firm’s profit maximization problem as defined above.
  3. 3.
    The government satisfies the balanced budget constraint by collecting income taxes, inheritance taxes, and social security payments, and distributing lump-sum transfers and pension benefits to individuals.
    $\underbrace{Tax^{\text{Labor}}}_{\text{Labor Income Tax}}+\underbrace{Tax^{\text{Asset}}}_{\text{Financial Income Tax}}+\underbrace{Tax^{\text{Bequest}}}_{\text{Inheritance Tax}}+\underbrace{SSC}_{\begin{subarray}{c}\text{Social Security}\\ \text{Contributions}\end{subarray}}$
    $=\underbrace{\int\Big{(}\underbrace{Tr}_{\begin{subarray}{c}\text{lump-sum}\\ \text{transfer}\end{subarray}}+\underbrace{ssb(e)\mathcal{I}_{p=1}}_{\text{Pension benefit}}\Big{)}F_{Y}(ds_{Y})}_{\text{between 25 and 79 years old}}+\underbrace{\int\Big{(}Tr+ssb(e)\mathcal{I}_{p=1}\Big{)}F_{O}(ds_{O})}_{\text{between 80 and 104 years old}}$
  4. 4.
    All the bequest is allocated equally to all the individuals, and this equation holds.
    $\int\Big{(}1-S(t)\Big{)}b(a^{\prime})\,F(ds_{Y})+\int\Big{(}1-S(t)\Big{)}b(a^{\prime})\,F(ds_{O})=B\bigg{(}\int\Big{(}S(t)\Big{)}\,F_{Y}(ds_{Y})+\int\Big{(}S(t)\Big{)}\,F_{O}(ds_{O})\bigg{)}$
  5. 5.
    Both asset and labor markets are cleared.
    • Asset Market Clearing Condition:
      $K^{D}=\int a^{\prime}_{Y}(s_{Y})F_{Y}(ds_{Y})+\int a^{\prime}_{O}(s_{O})F_{O}(ds_{O})$
    • Labor Market Clearing Condition: for each $j\in\{NL,L\}$,
[…]
  6. 6.
    The probability measures are consistent with the agent’s optimal choices, and therefore, these equations hold.
    $\forall B_{Y}\in\mathcal{B}(\textbf{S}_{Y}),F_{Y}(B_{Y})$ $=\frac{\int_{(a^{\prime},\min\{x_{-1}+1,n_{j}\},j,\phi,\eta_{1},t+1,e)\in B_{Y}}\mathbbm{1}\Big{\{}n(s_{Y})=W\Big{\}}\mathbbm{1}\Big{\{}j(s_{Y})=j_{-1}\Big{\}}S(t)F_{Y}(ds_{Y})}{\int_{s_{Y}\in\textbf{S}_{Y}}S(t)F(ds_{Y})}$
    $+\frac{\int_{(a^{\prime},1,j,\phi,\eta_{1},t+1,e)\in B_{Y}}\mathbbm{1}\Big{\{}n(s_{Y})=W\Big{\}}\mathbbm{1}\Big{\{}j(s_{Y})\neq j_{-1}\Big{\}}S(t)F_{Y}(ds_{Y})}{\int_{s_{Y}\in\textbf{S}_{Y}}S(t)F(ds_{Y})}$
    $+\frac{\int_{(a^{\prime},0,NW,\phi,\eta_{1},t+1,e)\in B_{Y}}\mathbbm{1}\Big{\{}n(s_{Y})=NW\Big{\}}S(t)F_{Y}(ds_{Y})}{\int_{s_{Y}\in\textbf{S}_{Y}}S(t)F(ds_{Y})}$
    $\forall B_{O}\in\mathcal{B}(\textbf{S}_{O}),F_{O}(B_{O})$ $=\frac{\int_{(a^{\prime},t+1,e)\in B_{O}}S(t)F_{O}(ds_{O})}{\int_{s_{O}\in\textbf{S}_{O}}S(t)F_{O}(ds_{O})}$
4 Calibration
4.1 Externally Set Parameters
Calibration is performed to align the model with observed data from Japan and the corresponding calibrated moments. Certain parameters are externally calibrated to capture agents’ retirement decisions, drawing on Japan’s economic institutions and relevant literature.
The capital depreciation rate is set at 8.8%. To determine the threshold between full-time and part-time work, $\mathcal{F}$, I analyze the distribution of hourly wages across working hours, noting sharp wage increase in wages beyond 35 hours per week. An upper bound on working hours, $\bar{h}$, of 105 hours per week.
The tax and social security system reflects Japan’s economic institutions from 2015 to 2019. National pension benefits and tax rates are based on regulations: the base pension benefit, $\underline{b}$, is set at 65,008 JPY per month, and the financial tax rate, $\tau_{r}$, is 20.315%. I also incorporate Japan’s progressive labor income tax and social insurance systems, including pensions and health insurance, despite their complexity. Agents begin receiving pensions at age sixty with a probability of 0.1372, while others start at age sixty-five. Pension eligibility is determined at age sixty, and the probabilities are independent of all other factors.
For other parameters, I adopt values from existing literature. The capital share of income, $\alpha$, and the elasticity of substitution between nonlinear and linear occupations, $\psi$, are set to 0.36 and 0.67, respectively, following Jang and Yum [2022]. For the Frisch elasticity, $\gamma$, I use the estimate of 1.50 from Keane [2022] after reviewing several studies on the parameter.
I standardize prices using the mean hourly wage and adjust the total factor productivity of the representative firm so that the model’s average hourly wage equals 1.
4.2 Calibration Result
Table 2 summarizes the internally calibrated parameters, computed to match the target moments in the data.
Figure 6 shows the share of workers in each occupation by age. As labor disutility increases independently of occupation, the model predicts a slight decline in the share of nonlinear occupations, while successfully replicating changes in the share of linear occupations. Figure 7 also reflects a similar trend between the data and the model. The presence of $\kappa$ motivates workers to reduce working hours after age 60, with the decline plateauing around age 70. At this point, workers are motivated to work more than 20 hours per week to maintain or increase pension benefits. Working fewer hours results in slightly lower pension benefits, while full-time work subjects them to the earnings test,explaining the preference for part-time work among older individuals. In Appendix, I compare graphs of wage, labor force participation, and asset holdings from the baseline model with the data.
Table 2: Internally Calibrated Parameters
Parameter Target Statistics
Value Description Data Model Description Data Source
$\beta$ 1.0094 Discount Factor 0.0107 0.01082 Real interest rate International Monetary Fund [2017]
$\mu_{\phi}$ 10.40 Mean of working hours 0.4226 0.4223 $\mathbb{E}[h|25\leq t\leq 59]$ JPSED
$\delta_{\phi}$ 1.900 S.D. of labor disutility 0.2665 0.2273 sd(log(h)) (25-59 years old) JPSED
$\kappa$ 0.04850 Coefficient in $\Phi$ 44.31 44.9187 LFP rate between 60 and 79 JPSED
$\xi$ 0.3900 Labor Force participation cost 5.6034 5.6090 Proportion of workers with $h<20$ per week (%) JPSED
$\nu$ 0.5570 Weight of NL laborforce 1.1605 1.1655 $\mathbb{E}[wmh^{\theta}gzf|j=NL]/\mathbb{E}[wmh^{\theta}gzf|j=L]$ JPSED
$\sigma_{\eta_{1}}$ 0.1600 Variance of $\eta_{1}$ 0.8884 0.8950 Share of all workers in NL JPSED
$\theta_{1}$ 0.4088 Curvature of $g_{NL}(\cdot)$ 0.1408 0.1514 Part-time penalty (NL) JPSED
$\theta_{2}$ 0.2480 Curvature of $g_{L}(\cdot)$ 0.07766 0.07740 Part-time penalty (L) JPSED
$\Omega_{1}$ 0.02130 Coefficient in $z_{N}L(\cdot)$ 0.3868 0.3931 Experience Premium in NL JPSED
$\Omega_{2}$ 0.01900 Coefficient in $z_{L}(\cdot)$3018 0.3026 Experience Premium in L JPSED
$\pi_{1}$0.1130 Coefficient in $\log f_{N}L(\cdot)$ 0.7564 0.7763 Wage reduction after sixty in NL JPSED
$\pi_{2}$0.06200 Coefficient in $\log f_{L}(\cdot)$7922 0.8082 Wage reduction after sixty in L JPSED
$\mu_{1}$43.00 Concern about leaving bequests 0.01283 0.020216 Inheritance rate JPSED
$\mu_{2}$ 1.700 Bequests as luxury goods 1.2753 1.1254 30th pct of Assets (80-105 years old) JHPS/KHPS
$\rho$ 0.3310 Coefficient in ssb(M) 0.1520 0.1519 $\mathbb{E}[ssb(M)]$ Ministry of Health, Labour and Welfare
  • 1
    The efficiency wage of each worker is $\frac{w_{j}m_{j}(\eta_{1})g_{j}(h)z_{j}(x,t)f_{j}(t)}{h}=w_{j}m_{j}(\eta_{1})h^{\theta_{j}}z_{j}(x,t)f_{j}(t)$.
  • 2
    Part-time penalty in occupation j is defined by $\mathbb{E}[wmh^{\theta}zf|j,h\geq\mathcal{F}]-\mathbb{E}[wmh^{\theta}zf|j=NL,h<\mathcal{F}]$. The difference is used as a measure of part-time penalty instead of ratios because it is a better way to gauge the curvature of the function, $g_{j}(\cdot)$.
  • 3
    Experience Premium in occupation j is also defined by $\mathbb{E}[wmh^{\theta}zf|j,50\leq t\leq 59]-\mathbb{E}[wmh^{\theta}zf|j,25\leq t\leq 34]$. For the same reason as part-time penalty, I use the difference, not the ratio.
  • 4
    Wage reduction after 60 years old in occupation j denotes $\mathbb{E}[wmh^{\theta}zf|j=NL,60\leq t\leq 69]/\mathbb{E}[wmh^{\theta}zf|j=NL,50\leq t\leq 59]$, which is the ratio of efficiency wage of workers between 60 and 69 years old to that of those between 50 and 59 years old.
  • 5
    LFP rate is an acronym for ”Labor force participation rate”.
Figure 6: Unconditional Occupation Share in Japan (Model vs Data)
Figure 7: Working Hours per week in Japan (Model vs Data)
4.3 Source of nonlinearity
The parameters driving nonlinearity in the model are analyzed, focusing on the part-time penalty, experience premium, and wage reduction after retirement. Nonlinearity is characterized by three indicators: the wage gap between full-time and part-time workers, the wage gap between workers in their fifties and sixties, and the wage ratio between workers in their seventies and sixties.
The analysis reduces the values of $\theta$, $\Omega$, and $\pi$ by half and examines the outcomes in partial equilibrium (Table 3). A lower $\theta$ encourages part-time work, increasing both the extensive and intensive labor margins for older workers, particularly in nonlinear occupations, and boosting capital supply. In contrast, $\Omega$ has minimal impact on nonlinearity, while $\pi$ partially affects the part-time penalty.
Table 3: Source of nonlinearity (Partial equilibrium)
Baseline $\theta\downarrow 50\%$ $\Omega\downarrow 50\%$ $\pi\downarrow 50\%$
Supply side
$\Delta$ Labor Supply(NL)(%) +8.649 -14.61 +1.082
$\Delta$ Labor Supply(L)(%) +7.101 -15.94 +1.604
$\Delta$ Labor Supply(Age:60-79)(%) +61.55 -30.80 +25.13
$\Delta$ Capital Supply(%) -7.034 -10.13 -2.766
$\Delta$ Tax Revenue(%) +5.019 -15.28 0.5190
Labor Market Indicators
LFP Rate(Age:60-79)(%) 44.9187 77.04 37.09 53.46
Part-time Rate (NL)(%) 10.9677 48.17 13.36 12.65
Part-time Rate (L)(%) 31.6372 57.18 39.01 34.56
Average Working Hours(Age:25-79) 0.4223 0.3806 0.4192 0.4177
NL/L: Population ratio 0.8950 0.9681 0.8901 0.9035
NL/L: Wage ratio 1.1655 1.184 1.1605 1.1591
Part-time penalty (NL) 0.1514 0.03233 0.1314 0.09217
Part-time penalty (L) 0.07740 0.01602 0.05307 0.05151
Experience Premium (NL) 0.3931 0.4790 0.09224 0.3822
Experience Premium (L) 0.3026 0.3525 0.1028 0.3036
Wage reduction after 60(NL) 0.7763 0.7812 0.7946 0.8527
Wage reduction after 60(L) 0.8082 0.8480 0.8168 0.8561
1616footnotetext: NL/L:Population ratio is the ratio of the number of Nonlinear workers to that of Linear workers. Also, NL/L:Wage ratio means the ratio of the average efficiency wage of Nonlinear workers to that of Linear workers.
5 Counterfactual Experiment
Several counterfactual experiments are conducted to evaluate both conventional and unconventional policy reforms. Although conventional policies are likely to reduce welfare, unconventional policies, including income tax reforms, increase welfare, while increasing output.
Note that in all the experiments, the real interest rate is about 1% , while there is no population growth and no technological growth in the stationary equilibria: the equilibria are dynamically efficient.
First, as shown in Table 4, eliminating the earnings test positively impacts consumption equivalence: the short-term consumption equivalence (CEV)171717 The short-term CEV refers to the consumption equivalent variation for individuals aged 25 immediately following the policy reform, based on the distribution of agents in the baseline model. In contrast, the long-term CEV represents the consumption equivalent variation for individuals of the same age between the two stationary equilibria For example, CEV at age t, denoted as $CEV_{t}$, is defined as follows: Let $\{c_{s}^{0},h_{s}^{0},a_{s+1}^{0}\}_{s=t}^{T}$ and $\{c_{s}^{1},h_{s}^{1},a_{s+1}^{1}\}_{s=t}^{T}$ represent the household’s allocation in an equilibrium of the baseline model and the compared allocation, respectively. $u\Big{(}(1+CEV_{t})c_{t}^{0},h_{t}^{0}\Big{)}+\Big{(}1-S(t)\Big{)}\mu(a_{t+1}^{0})+\sum_{s=t+1}^{T}\beta^{s-t}S(s-1)\bigg{\{}u\Big{(}(1+CEV_{t})c_{s}^{0},h_{s}^{0}\Big{)}+\Big{(}1-S(s)\Big{)}\mu(a_{s+1}^{0})\bigg{\}}$ $=u\Big{(}c_{t}^{1},h_{t}^{1}\Big{)}+\Big{(}1-S(t)\Big{)}\mu(a_{t+1}^{1})+\sum_{s=t+1}^{T}\beta^{s-t}S(s-1)\bigg{\{}u\Big{(}c_{s}^{1},h_{s}^{1}\Big{)}+\Big{(}1-S(s)\Big{)}\mu(a_{s+1}^{1})\bigg{\}}$ and long-term CEV at 25 years old are $0.001061\%$ and $0.01940\%$, respectively. Across all generations, the short-term CEV remains positive but small, as shown in Figure 8. This reform does not affect the extensive margin of older individuals but increases the intensive margin by $2.539\%$. Additionally, it boosts aggregate output, labor supply, capital supply, and tax revenue, though these increases remain under $1\%$.
Second, extending the pension eligibility age by five years delays benefits from age 60 to 65 and from 65 to 70. This reform reduces welfare, particularly for those approaching retirement. While it increases labor supply among older workers in both nonlinear and linear occupations, it reduces capital supply and output. The expanded labor supply also causes younger and middle-aged individuals to work less, leading to wage declines that lower earnings and discourage savings in anticipation of extended working years, as shown in Appendix. This result arises from capital adjustments driven by price changes and the intertemporal substitution of labor, as workers anticipate longer working lives. When comparing factors, these forces significantly contribute to the observed changes, whereas competition between younger and older individuals for labor supply has only a slight effect. Although Imrohoroğlu and Kitao [2012] demonstrates that extending the normal retirement age increases capital as individuals save to smooth consumption during periods without pension eligibility, this model allows workers to adjust their extensive margin to maintain consumption levels, which consequently reduces the capital supply.
Third, reducing employees’ pension benefits by half, resulting in a $28.66\%$ average decrease in pension benefits, lowers welfare in both the short and long term. Output declines by $0.0703\%$, and the increased labor supply from older individuals suppresses wages, discouraging labor supply and earnings for the working-age population. Workers in linear occupations, earning lower wages and holding fewer assets, experience a stronger income effect, increasing their labor supply in linear occupations more than in nonlinear ones.
Next, I examine the effects of unconventional policy reforms on welfare, output, and labor. Reducing $\theta_{NL}$ to the same level as $\theta_{L}$ boosts older individuals’ labor supply by $19.47\%$ and raises the LFP rate for those in their 60s and 70s by $10.64\%$. This increase in nonlinear labor supply also enhances capital supply and output, yielding a long-term CEV of $0.8773\%$.
Moreover, I cut $\pi_{1}$ and $\pi_{2}$ by half, corresponding to the age-penalty in nonlinear and linear occupations, respectively. This experiment increases labor supply, capital supply, and output but results in a small negative CEV due to the extended working years required.
Furthermore, enhancing tax credits by 1.5 times, including deductions like the basic and dependents’ deductions, lowers the marginal tax rate for workers. This reform boosts older workers’ labor participation by $5.35\%$ and increases hourly wages, encouraging continued employment. While tax revenue declines by $2.701\%$, the labor force expands across occupations, driving savings and increasing output by $2.207\%$.
Lastly, exempting pensions from income tax increases disposable income, boosting labor supply by $8.958\%$ and raising the labor force participation rate by $6.92\%$. This reform enhances welfare and increases output by $0.7708\%$.
These experiments show that conventional policies like extending pension eligibility and cutting benefits reduce welfare (CEV) and face resistance from middle-aged and older individuals, delaying implementation. In contrast, unconventional policies boost welfare and encourage increased labor supply, including greater participation in nonlinear occupations.
Table 4: Policy Experiments
Baseline Eliminate ETest Extend Pen Age by five years Lower Pension $\theta_{NL}\downarrow$ until $\theta_{L}$ $\pi\downarrow$ until 0.5 $\pi$ Increase Tax Credit No Tax on Pension
Consumption Equivalence
Short term CEV(Age:25)(%) +0.001061 -4.152 -1.880 -0.008137 -0.004552 +2.062 +0.2610
Long term CEV(Age:25)(%) +0.01940 -0.05688 -4.441 +0.8773 -0.09088 + 0.896 + 0.9529
Aggregate Change
$\Delta$Output(%) +0.3556 -0.0811 -0.0703 +2.247 +0.5620 +2.207 +0.7708
$\Delta$ Labor Supply(NL)(%) +0.04820 +1.400 -1.1090 +5.504 +0.1957 +1.763 +0.4959
$\Delta$ Labor Supply(L)(%) +0.03589 +1.023 + 1.069 -0.4785 +0.9715 +1.785 +0.9071
$\Delta$ Labor Supply(Age:60-79)(%) +2.539 +68.96 +48.92 +19.47 +9.324 +6.010 +8.958
$\Delta$ Capital Supply(%) +0.6505 -2.366 +0.02090 +1.442 +0.5843 +2.982 +0.9258
$\Delta$ Tax Revenue(%) +0.1932 -1.256 -4.241 +1.574 +1.574 -2.701 -0.1220
Labor Market Indicators
Real interest rate (%) 1.082 1.082 1.201 1.091 1.185 1.091 1.020 1.081
LFP Rate(Age:60-79)(%) 44.92 44.78 62.55 79.63 55.56 48.33 49.27 51.84
Part-time Rate (NL)(%) 10.97 9.800 24.24 31.17 41.69 10.40 11.98 11.89
Part-time Rate (L)(%) 31.64 31.25 39.10 50.23 32.11 33.56 34.69 35.93
AVG Working Hours(Age:25-79) 0.4223 0.4225 0.4009 0.4029 0.4039 0.4213 0.4258 0.4218
NL/L: Population ratio 0.8950 0.8929 0.9314 0.8992 1.018 0.8924 0.8718 0.8811
NL/L: Wage ratio 1.1655 1.166 1.152 1.154 1.198 1.164 1.167 1.165
Part-time penalty (NL) 0.1514 0.1303 0.02576 0.1457 0.03690 0.1231 0.1531 0.1501
Part-time penalty (L) 0.07740 0.07689 0.04881 0.06888 0.07138 0.06779 0.08553 0.08731
Experience Premium (NL) 0.3931 0.3785 0.3552 0.3746 0.4262 0.3734 0.4025 0.3791
Experience Premium (L) 0.3026 0.3050 0.2878 0.2942 0.3043 0.3007 0.3128 0.3010
Wage reduction after 60(NL) 0.7673 0.7960 0.9192 0.7212 0.7787 .8121 0.7645 0.7668
Wage reduction after 60(L) 0.8082 0.8095 0.9094 0.8057 0.8173 0.8258 0.7988 0.8003
  • 1
    Apart from the short-term CEV, all indicators reflect either the changes in the stationary equilibrium relative to the baseline model
    or the levels within the stationary equilibrium achieved in the experiment.
  • 2
    ”Eliminate ETest” denotes eliminating earnings test.
  • 3
    ”Extend Pend Age by five years” means extending pension eligibility age by five years.
Figure 8: Short-term CEV by Age(Conventional Policies)
1717footnotetext: The CEV for the ’Extend Pension Age’ policy is plotted only up to age 64 because individuals aged 65 to 69 lose pension eligibility under this reform, which is embedded within the value functions.
Figure 9: Short-term CEV by Age(Unconventional Policies)
6 Conclusion
This study reveals that workers in nonlinear occupations tend to retire earlier than those in linear occupations, based on analyses using JPSED and JHPS/KHPS data. This disparity arises from the high part-time penalties in nonlinear occupations, further magnified by the presence of pensions and assets, which elevate the reservation wage for workers in these occupations.
It also demonstrates that while reducing pension benefits decreases labor supply in nonlinear occupations, other policies effectively increase both the intensive and extensive margins of older workers. These heterogeneous effects arise from the interplay of income and substitution effects across different occupational types. When making policy plans, it is important to consider these heterogeneous effects.
For future research, it is essential to investigate additional factors influencing age-friendliness across occupations, such as working conditions, physical strain, and firms’ profit-maximizing strategies. A more comprehensive assessment of these aspects could guide the creation of more effective and targeted policies.
Moreover, examining the labor supply of older females has become increasingly important. While 43.72% of females aged 25–59 are employed part-time, their labor force participation rate remains at approximately 80% within the working-age population. As female workers continue to constitute a growing share of the labor market and approach retirement, understanding their retirement behavior alongside that of males is essential for developing inclusive and comprehensive labor market policies.
Appendix
A.1. Data Description in Japan
Figure 10: Unconditional Proportion of Workers’ Choices After Quitting Jobs in Japan: Male, 2015-2019
1818footnotetext: ”Leave” represents the proportion of workers who exit the labor force after leaving their occupation. ”Switch across” refers to transitions between different occupation types; for example, a nonlinear worker moving to a linear occupation. ”Switch within” indicates job changes within the same occupation category; for instance, a nonlinear worker moving to another nonlinear job.
Figure 11: Lifetime hourly wage (nonlinear vs. linear) in Japan: Male, 2015-2019
Figure 12: Change of Conditional Occupation Share over Age(nonlinear vs. linear) in Japan: Male and Female, 2009-2019
A.2. Data Description in US
Figure 13: Hourly Wage Change over Working Hours(nonlinear vs. linear) in US: Male, 2009-2019
Figure 14: Change of Conditional Occupation Share over Age(nonlinear vs. linear) in US: Male, 2009-2019
Figure 15: Lifetime hourly wage (nonlinear vs. linear) in US: Male, 2009-2019
A.3. Cross-country Data Description
Figure 16: Labor force participation rate, cross-country comparison: Male, 2015-2019
A.4. Calibration result
Figure 17: Wage difference in Japan (Model vs Data)
Figure 18: LFP rate in Japan (Model vs Data)
Figure 19: Asset in Japan (Model vs Data)
A.5. Source of nonlinearity
Figure 20: Unconditional Occupation Share (Source of nonlinearity)
Figure 21: Wage difference (Source of nonlinearity)
Figure 22: Effective Labor Supply (Source of nonlinearity)
Figure 23: Working Hours per week (Source of nonlinearity)
Figure 24: LFP (Source of nonlinearity)
Figure 25: Earnings (Source of nonlinearity)
Figure 26: Asset (Source of nonlinearity)
A.6. Counterfactual Experiment(Conventional Policy)
Figure 27: Unconditional Occupation Share(Conventional Policy)
Figure 28: Wage difference (Conventional Policy)
Figure 29: Effective Labor Supply(Conventional Policies)
Figure 30: Weekly Working Hours (Conventional Policy)
Figure 31: LFP rate (Conventional Policy)
Figure 32: Earnings (Conventional Policy)
Figure 33: Asset (Conventional Policy)
A.7. Counterfactual Experiment(Unconventional Policy)
Figure 34: Unconditional Occupation Share (Unconventional Policy)
Figure 35: Wage difference (Unconventional Policy)
1818footnotetext: When I reduce $\theta_{NL}$ to $\theta_{L}$, since both of $\frac{K}{L}$ and $\frac{EL}{E_{1}}{L_{1}}$ decrease, leading to a reduction in $w_{1}$, which represents the payment per unit of efficiency labor in nonlinear occupations. However, nonlinear workers with high labor disutility increase their productivity and receive a higher hourly wage, while those with lower labor disutility reduce their working hours due to the weakened nonlinearity. As a result, the overall wage level for a nonlinear worker increases.
Figure 36: Effective Labor Supply(Unconventional Policy)
Figure 37: Working Hours per week (Unconventional Policy)
Figure 38: LFP (Unconventional Policy)
Figure 39: Earnings (Unconventional Policy)
Figure 40: Asset (Unconventional Policy)